OpenStudy (anonymous):

Given the linear equation y = -6x – 2, find the y-coordinates of the points (-1, ), (-2, ), and (1, ). Please show all of your work. Plot these points and graph the linear equation

OpenStudy (anonymous):

i suck at this stuff

OpenStudy (kainui):

So for the first one you're given -1 for your x, just plug that in to the formula like this: y=(-6*-1)-2 and you get y=4, so the coordinate is (-1,4) No one's born knowing how to do this kind of stuff, but anyone can learn if they decide they want to. =)

OpenStudy (anonymous):

ok how do i get the other two

OpenStudy (kainui):

Just repeat what I showed you, but with the other x-coordinates.

OpenStudy (anonymous):

so would it be like this x=(-6x 1)

OpenStudy (phi):

they give you y = -6x – 2 that is short hand for y is -6 times x , then subtract 2 if they tell you x= -1 do this y = (-6*-1) -2 then do the arithmetic

OpenStudy (anonymous):

y=(-6*-2)-2

OpenStudy (phi):

yes. now do -6 * -2 = ?

OpenStudy (anonymous):

12

OpenStudy (phi):

so y= 12-2 finish it

OpenStudy (anonymous):

12-2=10

OpenStudy (phi):

yes so y=10 that is the 2nd answer last one, replace x with 1

OpenStudy (anonymous):

Coordinate is(-2,10)

OpenStudy (anonymous):

Y=(-6*-1)-2 -6*-1=6 6-2=4 Coordinate is (-1,4) Y=(-6*-2)-2 -6*-2=12 12-2=10 Coordinate is(-2,10) -6*1=-6 -6-2=-8 Coordinate is (1,-8)

OpenStudy (anonymous):

rght

OpenStudy (phi):

yes looks good.

OpenStudy (anonymous):

thank you

OpenStudy (phi):

they still want you to plot the points. can you do that?

OpenStudy (anonymous):

ya i'm good with that

OpenStudy (phi):

you know it's right if all 3 points are on the same line.

OpenStudy (anonymous):

ya